OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, 7) and (–4, –8)? A. –3 B. -1/3 C. 1/3 D. 3

OpenStudy (uri):

use the slope formula

pooja195 (pooja195):

\[\huge~\frac{ y2-y2 }{ x2-x1 }\]

pooja195 (pooja195):

(x1,y1) (x2,y2)

OpenStudy (uri):

omg why am i helping?! not supposed to help, byez

OpenStudy (anonymous):

why not D: @uri

OpenStudy (anonymous):

and i understand the formula i did it and got -4/6 i just am lossed from there.

OpenStudy (black_ninja123):

its actually D

OpenStudy (anonymous):

how do you figure?

OpenStudy (black_ninja123):

\[\frac{ -8-7 }{ -4-1 }\]

OpenStudy (uri):

i love you and all but i cant bear kids like pooja @Swordsman1305

OpenStudy (black_ninja123):

\[\frac{ -15 }{ -5 }=3\]

OpenStudy (anonymous):

oh.. well pooj left

OpenStudy (black_ninja123):

pooja formula was wrong she wrote y2-y2 and its actually y2-y1

OpenStudy (anonymous):

black i thought i was supposed to do 8-4 and 7-1?

OpenStudy (anonymous):

yes.. i think so>? pooja right?

OpenStudy (black_ninja123):

y2 is -8 and y1 is 7 so -8-7= -15

OpenStudy (black_ninja123):

x2 is -4 and x1 is 1 -4-1=5

OpenStudy (anonymous):

well anyway i dont do y2= -8 y1= -4 x2 = 7 x1= 1 and then -8-(-4) 7-1 and get -4/6?

OpenStudy (black_ninja123):

-5*

OpenStudy (anonymous):

so what i did is the wrong way?

OpenStudy (black_ninja123):

y1 is 7 not -4

OpenStudy (black_ninja123):

x1=1 y1=7 x2=-4 y2=-8

OpenStudy (anonymous):

ugh this is soo confusing i went off what you said here http://openstudy.com/study#/updates/556de732e4b01de5673c84b4 @Black_ninja123

OpenStudy (black_ninja123):

Oh i'm so sorry for that i got the numbers mixed up

OpenStudy (anonymous):

so it was the wrong answer?

OpenStudy (black_ninja123):

Use the example i just gave you

OpenStudy (black_ninja123):

the answer is right

OpenStudy (black_ninja123):

use this calc http://www.mathwarehouse.com/calculators/slope-formula-calculator.php

OpenStudy (anonymous):

so the first one is right.. you just mixed up the formula set up a bit?

OpenStudy (anonymous):

ok well just to clarrify this one is D and the last one i asked was the right one that you gave? correct?

OpenStudy (anonymous):

ok thanks i gave u medal

OpenStudy (black_ninja123):

yes sorry I made things confusing

OpenStudy (black_ninja123):

I a horrible at explaining things

OpenStudy (black_ninja123):

am*

OpenStudy (anonymous):

its ok i tagged u to another question.. perhaps u want another medal? :D